OpenStudy (anonymous):

Divide, then simplify your answer. Please/Thanks 1/n ÷ n-4/4n A) n-4/4 B) 4/n C) 1(4n)/n(n-4) D) 4/n-4

OpenStudy (anonymous):

\[\large \frac{1}{n} \div \frac{n-4}{4n}=\frac{1}{n} \times \frac{4n}{n-4}\] \[\large =\frac{4}{n-4}\]

OpenStudy (anonymous):

Thank You!